The table shows a firm’s total and marginal costs.

Output Total cost ($) Marginal cost ($)

1 200 20

2 215 15

3 225 10

4 240 15

5 260 20

Calculate the Average Fixed Cost (AFC) of producing 6 units of output.


1
Expert's answer
2021-12-03T13:07:04-0500

When output is zero total cost would be:

Total cost (0 unit) = total cost at 1 unit - marginal cost at 1 unit

Total cost (0 unit) = 200 - 20 = 180

This total cost includes only fixed cost due to zero production. So fixed cost is 180.

AFC of producing 6 units = 180/6 = $30 per unit